OpenStudy (anonymous):

Determine the Molarity of an acid solution; if 500mL of Sulfuric acid (H2SO4)totally neutralizes 750mL of 6.412M Sodium Hydroxide (NaOH).

OpenStudy (anonymous):

can you find the number of moles of NaOH that that was neutralized?

OpenStudy (anonymous):

Can you walk me through it?

OpenStudy (anonymous):

use the formula \[\large c=\frac{ n }{ v }\] c=concentration (measured in M) n=amount of the substance/ no. of moles (measured in mol) v=volume of the substance (measured in L)

OpenStudy (anonymous):

\[\large n=cv\] \[\large n= 6.412M~\times~750\times 10^{-3}L\] \[\large n=??\]

OpenStudy (anonymous):

that is finding the number of moles of NaOH..

OpenStudy (anonymous):

what do you get for n??

OpenStudy (anonymous):

4.809

OpenStudy (anonymous):

sorry for not replying earlier....i had to do something for a friend...

OpenStudy (anonymous):

you got the number of moles right....do you know about mole ratio??

OpenStudy (anonymous):

@lovely_layne